OpenStudy (chaotic_butterflies):

Need help with systems of equations word problem! The research department of a company that manufactures children's fruit drinks is experimenting with a new flavor. a 17.5% fructose solution is needed but only 10%and 20% solutions are available. How many gallons of the 10% fructose solution should be mixed with the 20% fructose solution to obtain 20 gallons of 17.5% fructose solution? I know how to solve a systems of equations problem, but I've really struggled to make the word problems into the equations themselves.

OpenStudy (chaotic_butterflies):

@mathstudent55 when you're not busy, could you help me?

OpenStudy (chaotic_butterflies):

@myininaya

OpenStudy (chaotic_butterflies):

@Owlcoffee

OpenStudy (plasmataco):

yo I got u

OpenStudy (chaotic_butterflies):

Thank you so much already XD

OpenStudy (plasmataco):

I hope

OpenStudy (majesticgalaxy):

I'll help you out, this is a long one so I'll try to explain the problem as best as i can ;)

OpenStudy (majesticgalaxy):

Let the amount of 20% solution be x gallons Amount of 10% solution = 20 - x gallons 0.20 ( x) + 0.10 ( 20 - x) = 0.175 ( 20) 0.20x + 2 - 0.10x = 3.5 0.10x = 1.5 x = 1.5 / 0.10 = 15 gallons ANSWER = 15 gallons ( 20%) 5 gallons ( 10%) CHECK 0.20 ( 15) = 3 0.10 ( 5) = 0.5 Total 3.5 /20 X 100 = 17.5%

OpenStudy (plasmataco):

eziest way to do this is make a table

OpenStudy (plasmataco):

or listen to @MajesticGalaxy lel

OpenStudy (majesticgalaxy):

your answer would be 17.5%

OpenStudy (chaotic_butterflies):

if you're speaking of matrices my math teacher decided to skip that...

OpenStudy (plasmataco):

nah. Just a 3 by 3 table

OpenStudy (chaotic_butterflies):

isn't there supposed to be another variable?

OpenStudy (plasmataco):

nvm

OpenStudy (plasmataco):

hang on.

OpenStudy (majesticgalaxy):

No i do not think so

OpenStudy (chaotic_butterflies):

No, I know this wants me to do it a specific way, it's all about systems of equations

OpenStudy (plasmataco):

ok, so one of the equations is 0.1x+0.2y=20*0.175

OpenStudy (plasmataco):

the other I think is x+y=20

OpenStudy (chaotic_butterflies):

y being the 20% fructose?

OpenStudy (plasmataco):

yes.

OpenStudy (chaotic_butterflies):

THANK YOU

OpenStudy (plasmataco):

No prob, make sure u check if it makes sense!

OpenStudy (chaotic_butterflies):

I'll solve and see what I get

OpenStudy (plasmataco):

I'll try 2

OpenStudy (plasmataco):

K, I got x=5, y=15

OpenStudy (chaotic_butterflies):

I got that too! And it makes sense! Thank you so very much!

OpenStudy (plasmataco):

no prob!

OpenStudy (mathstudent55):

If you look at @MajesticGalaxy 's solution, you will see that he is very close to the system of equations you need. He skipped defining variable y, and used substitution. Below is a copy of his solution. I will add what you need to have a system of solutions in between lines of *******. Let the amount of 20% solution be x gallons ********************************* Amount of 10% solution is y You need a total of 20 gallons, so x + y = 20 Solving for y, the amount of 10% solution, you get ********************************* Amount of 10% solution = y = 20 - x gallons ********************************** The amount of fructose is 0.20x + 0.10y = 0.175 * 20 Since we know that y = 20 - x, we substitute to get: ********************************** 0.20 ( x) + 0.10 ( 20 - x) = 0.175 ( 20) 0.20x + 2 - 0.10x = 3.5 0.10x = 1.5 x = 1.5 / 0.10 = 15 gallons ANSWER = 15 gallons ( 20%) 5 gallons ( 10%) CHECK 0.20 ( 15) = 3 0.10 ( 5) = 0.5 Total 3.5 /20 X 100 = 17.5%